OpenStudy (anonymous):

Accumulated value after investing $4500.00 for 8 years at 6% quarterly

OpenStudy (anonymous):

\[4500\times \left(1+\frac{.06}{4}\right)^{32}\] and a calculator

OpenStudy (anonymous):

the 32 because \(8\times 4=32\)

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

yw